# Convergence and error control of consistent PINNs for elliptic PDEs

### v1

@@ 6.2 Discrete $H^{1/2}(\partial\Omega)$ norms
In this section, we introduce discrete $H^{1/2}(\partial\Omega)$ norms and discuss their accuracy in computing the true $H^{1/2}(\partial\Omega)$ norm for functions $g$ in our model class ${\cal G}$. The discrete norm we introduce will only uses the values of $g$ at the data sites and can therefore be computed from the given data. We fix any $k\geq 1$ and let ${\cal Z}:=\overline{G}_{k,r}=\{{\bf z}_{j}\}_{j=1}^{\bar{m}}=G_{k,r}\cap\partial\Omega$, see (3.16), be the grid points on the boundary $\partial\Omega$. The number of grid points in ${\cal Z}$ is $\overline{m}=2d[r2^{k}]^{d-1}$.
@@
$R_{m}^{*}({\cal G})_{H^{1/2}(\partial\Omega)}\asymp m^{-\beta},\quad m\geq 1,\quad\beta=\frac{\bar{s}-1}{d-1}.$ (6.7)
@@
$\|g-\overline{S}_{k}(g)\|_{H^{1/2}(\partial\Omega)}\leq C\|g\|_{Tr(\overline{\cal B})}\bar{m}^{-\beta},$ (6.10)
@@
$\left|\|g\|_{H^{1/2}(\partial\Omega)}-\|\overline{S}_{k}(g)\|_{H^{1/2}(\partial\Omega)}\right|\leq\|g-\overline{S}_{k}(g)\|_{H^{1/2}(\partial\Omega)}\leq C\|g\|_{Tr(\overline{\cal B})}\bar{m}^{-\beta}.$
We concentrate therefore on finding a discrete $H^{1/2}(\partial\Omega)$ norm for the functions $\overline{S}_{k}(g)$, $g\in{\cal G}$. Note that $\overline{S}_{k}(g)$ is completely determined by the values of $g$ at the data sites $\overline{G}_{k,r}$.
Given $k\geq 0$, let ${\cal T}_{k}:={\cal T}_{k}(\partial\Omega)$ denote the collection of all $d-1$ dimensional simplicies given by the Kuhn-Tucker partition of the dyadic cubes $I\in{\cal D}_{k}(\partial\Omega)$ of the boundary $\partial\Omega$. For each dyadic cube $I\in{\cal D}_{k}(\partial\Omega)$, there are $(d-1)!$ simplices contained in $I$
@@
and prove that the the two semi-norms $|\cdot|_{H^{1/2}(\partial\Omega)}^{*}$ and $|\cdot|_{H^{1/2}(\partial\Omega)}$ are equivalent on $V^{r}({\cal T}_{k})$.
We begin with the following lemma.
@@
We fix $d,\,r$, and a pair $T,T^{\prime}\in{\cal T}_{k}:={\cal T}_{k}(\partial\Omega)$. Let us first consider the case when $k\leq 4$. Let ${\cal N}(S):={\cal N}_{T,T^{\prime}}(S)$ denote the expression in (6.12) involving the sum, and ${\cal N}^{\prime}(S)$ the expression involving the integral. Both are semi-norms on the finite dimensional linear space $X=X(T,T^{\prime})$ of functions from $V^{r}({\cal T}_{k})$ restricted to $T\cup T^{\prime}$. A function $S\in X$ satisfies ${\cal N}(S)=0$ if and only if $S$ is constant on $T$, $T^{\prime}$ and, additionally, $S$ is constant on $T\cup T^{\prime}$ when $T$ and $T^{\prime}$ touch. A similar statement holds for the semi-norm ${\cal N}^{\prime}$. It follows that for any pair $T,T^{\prime}$, we have
@@
We now consider the case $k>4$. Given a pair $T,T^{\prime}$, we let $c(T,T^{\prime})$ be the largest number such that the lower inequality in (6.13) holds uniformly for all $S\in X(T,T^{\prime})$. Similarly $C(T,T^{\prime})$ is the smallest constant so that the upper inequality of (6.13) holds uniformly on $X(T,T^{\prime})$. The above argument using equivalence of semi-norms shows that there always exist positive constants $c(T,T^{\prime}),C(T,T^{\prime})$. We are left to show that there are constants $0<c^{*}<C^{*}<\infty$, depending only on $d$ and $r$, for which $c(T,T^{\prime})\geq c^{*}$ and $C(T,T^{\prime})\leq C^{*}$ holds for all pairs $T,T^{\prime}$ such that $T,T^{\prime}\in{\cal T}_{k}$ and $k>4$.
We consider two cases. Case 1: $T\cap T^{\prime}\neq\emptyset$. In this case there is a pair $T_{0},T_{0}^{\prime}\in{\cal T}_{4}$, and a linear mapping consisting of a translation and a dilation with factor $2^{-k+4}$ that rigidly transforms $T_{0},T_{0}^{\prime}$ to the pair $T,T^{\prime}$. If we use this linear transformation to change variables in (6.2) we obtain the validity of (6.2) for this case with the same constants $c_{1},C_{1}$.
Case 2: $T\cap T^{\prime}=\emptyset$. First note that from the $L_{2}$ discretization of the previous section, we have that there are constants $c_{2},C_{2}$ depending only on $r,d$ such that for each $S\in V^{r}({\cal T}_{k}(\partial\Omega))$
@@
and constants in $\lesssim$ depending only on $r,d$. Note that the optimal recovery rate of ${\cal G}$ in the $H^{1/2}(\partial\Omega)$ is $\bar{m}^{-\beta}$.
@@
$|S|_{H^{1/2}(\partial\Omega)}^{*}=\left[\sum_{{\bf z}\not={\bf z}^{\prime}\in\overline{G}_{k,r}}\frac{1}{\overline{m}^{2}}\frac{|S({\bf z})-S({\bf z}^{\prime})|^{2}}{|{\bf z}-{\bf z}^{\prime}|^{d}}\right[^{1/2}\asymp\left[\sum_{T,T^{\prime}\in{\cal T}_{k}}\sum_{\begin{subarray}{c}{\bf z}\not={\bf z}^{\prime}\in\overline{G}_{k,r}\\ {\bf z}\in T,\ {\bf z}^{\prime}\in T^{\prime}\end{subarray}}\frac{1}{\overline{m}^{2}}\frac{|S({\bf z})-S({\bf z}^{\prime})|^{2}}{|{\bf z}-{\bf z}^{\prime}|^{d}}\right]^{1/2}.$
@@
In particular this holds for $S=\overline{S}_{k}(g)$ whenever $g\in{\cal G}$. If we combine (6.17) with (6.10) and observe that $|g|^{*}_{H^{1/2}(\partial\Omega)}=|\overline{S}_{k}(g)|^{*}_{H^{1/2}(\partial\Omega)}$ we obtain the theorem. ∎
@@
Arguing as we have done in §6.1, we have for $\alpha=\frac{\bar{s}-1/2}{d-1}>\beta$ and any $g\in Tr(\overline{\cal B})$
@@
For any $g\in{\cal G}=Tr(U(\overline{B}))$, where $\overline{B}=B^{\bar{s}}_{\infty}(L_{2}(\Omega)$ with $\bar{s}>d/2$, and any $\overline{m}\geq 1$, we have
@@ 7 A discrete loss function with error control
In this section, we introduce a discrete loss function ${\cal L}^{*}$ which is a surrogate for the theoretical loss function ${\cal L}_{T}$ of (5.1). The advantage of ${\cal L}^{*}$ is that it can be computed directly from the data $({\bf f},{\bf g})$. Let $k,r$ be fixed and let ${\cal X}=G_{k,r}\subset\overline{\Omega}$ and ${\cal Z}=\overline{G}_{\bar{k},r}$ be the data sites of $\partial\Omega$. Note that $\bar{k}$ could be different from $k$. This fixes $\widetilde{m}=\#(G_{k,r})$ and $\overline{m}:=\#(\overline{G}_{\bar{k},r})$ for the remainder of this section.
@@
In this and the following two sections, the results and the proofs of the results are simplest in the case $d\geq 3$. When $d=2$, the results and proofs are clouded by the appearance of log factors. For this reason, we state the results in correct form (including the log factors) for all cases $d\geq 2$ but give the exposition and proofs only for $d\geq 3$. The proofs for the case $d=2$ are given in the Appendix, §12.7.
Let $\|\cdot\|^{*}_{L_{\gamma}}$ be the discrete norm, see (6.3), introduced in §6.1 for the following choices of
$\gamma=\begin{cases}\frac{2d}{d+2},\quad\quad&d\geq 3,\\ \\ 1+[\log(\widetilde{m})]^{-1},\,\,\,\quad&d=2,\end{cases}$
and let $\|\cdot\|^{*}_{H^{1/2}}$ be the discrete norm, see (6.11), for $H^{1/2}(\partial\Omega)$ introduced in §6.2. Given the data vectors for $({\bf f},{\bf g})$, we introduce the discrete loss function
@@
which is defined whenever $v$ and $\Delta v$ are continuous on $\overline{\Omega}$. Notice that to compute ${\cal L}^{*}(v)$ we use only the values of $v$ at the data sites.
@@
It is natural that $\|v\|_{{\cal U}}$ enters into the error bounds since we need some control on $v$ away from the data. However, this means that when we try to obtain a good approximation to $u$ (in $H^{1}(\Omega)$), through optimization, then the norm $\|\cdot\|_{{\cal U}}$ will have to enter the picture [4].
@@
This error estimator can be used to monitor the error in any optimization scheme that attempts to minimize the loss ${\cal L}^{*}(v)$. For example, suppose that $\Sigma=\Sigma_{n}$ is a linear space of dimension $n$ or a nonlinear manifold determined by $n$ parameters such that each $S\in\Sigma$ is continuous and has a continuous Laplacian $\Delta S$. If at any stage of the optimization procedure we have an $S\in\Sigma_{n}$ then we can use (7.3) to bound the error of $\|u-S\|_{H^{1}(\Omega)}$. Of course, a good estimate requires a bound for $\|S\|_{\cal U}$.
@@
Note that for fixed values of $\widetilde{m}$ and $\overline{m}$, the losses ${\cal L}$ and ${\cal L}^{*}$ are equivalent up to a constant depending upon the parameters $\widetilde{m}$ and $\overline{m}$. This means that if the original loss ${\cal L}$ is driven to 00, i.e. we are perfectly interpolating the data, then both losses give the same error control. However, in the situation where the data is not interpolated, the new loss ${\cal L}^{*}$ gives control on the error according to Theorem 7.2, while the original PINNs loss ${\cal L}$ does not.

### v2

@@ 6.2 Discrete $H^{1/2}(\partial\Omega)$ norms
In this section, we introduce discrete $H^{1/2}(\partial\Omega)$ norms and discuss their accuracy in computing the true $H^{1/2}(\partial\Omega)$ norm for functions $g$ in our model class ${\cal G}$. The discrete norm we introduce only uses the values of $g$ at the data sites and can therefore be computed from the given data. We fix any $k\geq 1$ and let ${\cal Z}:=\overline{G}_{k,r}=\{{\bf z}_{j}\}_{j=1}^{\overline{m}}=G_{k,r}\cap\partial\Omega$, see (3.15), be the grid points on the boundary $\partial\Omega$. The number of grid points in ${\cal Z}$ is $\overline{m}=2d[r2^{k}]^{d-1}$.
@@
$R_{\overline{m}}^{*}({\cal G})_{H^{1/2}(\partial\Omega)}\asymp{\overline{m}}^{-\beta},\quad{\overline{m}}\geq 1,\quad\beta=\frac{\bar{s}-1}{d-1}.$ (6.7)
@@
$\|g-\overline{S}_{k}(g)\|_{H^{1/2}(\partial\Omega)}\leq C\|g\|_{Tr(\overline{\cal B})}\overline{m}^{-\beta},$ (6.10)
@@
$\left|\|g\|_{H^{1/2}(\partial\Omega)}-\|\overline{S}_{k}(g)\|_{H^{1/2}(\partial\Omega)}\right|\leq\|g-\overline{S}_{k}(g)\|_{H^{1/2}(\partial\Omega)}\leq C\|g\|_{Tr(\overline{\cal B})}\overline{m}^{-\beta}.$
We concentrate therefore on finding a discrete $H^{1/2}(\partial\Omega)$ norm for the functions $\overline{S}_{k}(g)$, $g\in{\cal G}$. Note that $\overline{S}_{k}(g)$ is completely determined by the values of $g$ at the data sites $\overline{G}_{k,r}$. For each $k\geq 0$, we define ${\cal D}_{k}(\partial\Omega):=\partial\Omega\cap{\cal D}_{k}(\overline{\Omega})$ to be the partition of $\partial\Omega$ into dyadic cubes $I$ of side length $2^{-k}$. Given $k\geq 0$, let ${\cal T}_{k}:={\cal T}_{k}(\partial\Omega)$ denote the collection of all $d-1$ dimensional simplicies given by the Kuhn-Tucker partition of the dyadic cubes $I\in{\cal D}_{k}(\partial\Omega)$ of the boundary $\partial\Omega$. For each dyadic cube $I\in{\cal D}_{k}(\partial\Omega)$, there are $(d-1)!$ simplices contained in $I$
@@
and prove that the the two semi-norms $|\cdot|_{H^{1/2}(\partial\Omega)}^{*}$ and $|\cdot|_{H^{1/2}(\partial\Omega)}$ are equivalent on $V^{r}({\cal T}_{k})$. We begin with the following lemma.
@@
We fix $d,\,r$, and a pair $T,T^{\prime}\in{\cal T}_{k}:={\cal T}_{k}(\partial\Omega)$. Let us first consider the case when $k\leq 4$. Let ${\cal N}(S):={\cal N}_{T,T^{\prime}}(S)$ denote the expression in (6.12) involving the sum, and ${\cal N}^{\prime}(S)$ the expression involving the integral. Both are semi-norms on the finite dimensional linear space $X=X(T,T^{\prime})$ of functions from $V^{r}({\cal T}_{k})$ restricted to $\overline{T}\cup\overline{T^{\prime}}$. A function $S\in X$ satisfies ${\cal N}(S)=0$ if and only if $S$ is constant on $\overline{T}$, $\overline{T^{\prime}}$ and, additionally, $S$ is constant on $\overline{T}\cup\overline{T^{\prime}}$ when $\overline{T}$ and $\overline{T^{\prime}}$ touch. A similar statement holds for the semi-norm ${\cal N}^{\prime}$. It follows that for any pair $T,T^{\prime}$, we have
@@
We now consider the case $k>4$. Given a pair $T,T^{\prime}$, we let $c(T,T^{\prime})$ be the largest number such that the lower inequality in (6.13) holds uniformly for all $S\in X(T,T^{\prime})$. Similarly $C(T,T^{\prime})$ is the smallest constant so that the upper inequality of (6.13) holds uniformly on $X(T,T^{\prime})$. The above argument using equivalence of semi-norms shows that there always exist positive constants $c(T,T^{\prime}),C(T,T^{\prime})$. We are left to show that there are constants $0<c^{*}<C^{*}<\infty$, depending only on $d$ and $r$, for which $c(T,T^{\prime})\geq c^{*}$ and $C(T,T^{\prime})\leq C^{*}$ holds for all pairs $T,T^{\prime}$ such that $T,T^{\prime}\in{\cal T}_{k}$ and $k>4$. We consider two cases. Case 1: $\overline{T}\cap\overline{T^{\prime}}\neq\emptyset$. In this case, there is a pair $T_{0},T_{0}^{\prime}\in{\cal T}_{4}$, and a linear mapping consisting of a translation and a dilation with factor $2^{-k+4}$ that rigidly transforms $T_{0},T_{0}^{\prime}$ to the pair $T,T^{\prime}$. If we use this linear transformation to change variables in (6.2) we obtain the validity of (6.2) for this case with the same constants $c_{1},C_{1}$.
Case 2: $\overline{T}\cap\overline{T^{\prime}}=\emptyset$. First note that from the $L_{2}$ discretization of the previous section, we have that there are constants $c_{2},C_{2}$ depending only on $r,d$ such that for each $S\in V^{r}({\cal T}_{k}(\partial\Omega))$
@@
and constants in $\lesssim$ depending only on $r,d$. Note that the optimal recovery rate of ${\cal G}$ in the $H^{1/2}(\partial\Omega)$ is $\overline{m}^{-\beta}$.
@@
$|S|_{H^{1/2}(\partial\Omega)}^{*}=\left[\sum_{{\bf z}\not={\bf z}^{\prime}\in\overline{G}_{k,r}}\frac{1}{\overline{m}^{2}}\frac{|S({\bf z})-S({\bf z}^{\prime})|^{2}}{|{\bf z}-{\bf z}^{\prime}|^{d}}\right[^{1/2}\asymp\left[\sum_{T,T^{\prime}\in{\cal T}_{k}}\sum_{\begin{subarray}{c}{\bf z}\not={\bf z}^{\prime}\in\overline{G}_{k,r}\\ {\bf z}\in{\overline{T}},\ {\bf z}^{\prime}\in{\overline{T}^{\prime}}\end{subarray}}\frac{1}{\overline{m}^{2}}\frac{|S({\bf z})-S({\bf z}^{\prime})|^{2}}{|{\bf z}-{\bf z}^{\prime}|^{d}}\right]^{1/2}.$
@@
In particular this holds for $S=\overline{S}_{k}(g)$ whenever $g\in{\cal G}$. If we combine (6.17) with (6.10) and observe that $|g|^{*}_{H^{1/2}(\partial\Omega)}=|\overline{S}_{k}(g)|^{*}_{H^{1/2}(\partial\Omega)}$, we obtain the theorem. ∎
@@
Arguing as we have done in §6.1, we have for $\alpha=\frac{\bar{s}-1/2}{d-1}>\beta$ and any $g\in Tr(\overline{\cal B})$,
@@
For any $g\in{\cal G}=Tr(U(\overline{B}))$, where ¯$\overline{\cal B}=B^{\bar{s}}_{\infty}(L_{2}(\Omega))$ with $\bar{s}>d/2$, and any $\overline{m}\geq 1$, we have
@@ 7 A discrete loss function with error control
In section §4, we have introduced several numerical methods for finding a near optimal recovery to the solution $u\in{\cal U}_{\rm data}$ from the given data. In this section, we analyse an alternative approach that is based on optimization (such as PINNs). We consider a discrete loss function ${\cal L}^{*}$ which is a surrogate for the theoretical loss function ${\cal L}_{T}$, see (1.5). The advantage of ${\cal L}^{*}$ is that it can be computed directly from the data $({\bf f},{\bf g})$. Let $k,r$ be fixed and let ${\cal X}=G_{k,r}\subset\overline{\Omega}$ and ${\cal Z}=\overline{G}_{\bar{k},r}$ be the data sites of $\partial\Omega$. Note that $\bar{k}$ could be different from $k$. This fixes $\widetilde{m}=\#(G_{k,r})$ and $\overline{m}:=\#(\overline{G}_{\bar{k},r})$ for the remainder of this section.
@@
In this and the following two sections, the results and their proofs are simplest in the case $d\geq 3$. When $d=2$, log factors appear. For this reason, we state the results in correct form (including the log factors) for all cases $d\geq 2$, but give the proofs only for $d\geq 3$. The proofs for the case $d=2$ are given in the Appendix, §12.7.
Let $\|\cdot\|^{*}_{L_{\gamma}}$ be the discrete norm, see (6.3), introduced in §6.1 for the following choices of $\gamma$,
$\gamma=\begin{cases}\frac{2d}{d+2},\quad\quad&d\geq 3,\\ \\ 1+[\log(\widetilde{m})]^{-1},\,\,\,\quad&d=2,\end{cases}$ (7.1)
and let $\|\cdot\|^{*}_{H^{1/2}}$ be the discrete $H^{1/2}(\partial\Omega)$ introduced in §6.2, see (6.11), (6.18), and (6.20). Given the data vectors $({\bf f},{\bf g})$, we introduce the discrete loss function
@@
which is defined whenever $v$ and $\Delta v$ are continuous on $\overline{\Omega}$.
@@
It is natural that $\|v\|_{{\cal U}}$ enters into the error bounds since we need some control on $v$ away from the data. This means that when we try to obtain a good approximation to $u$ (in $H^{1}(\Omega)$) through optimization, then either the norm $\|\cdot\|_{{\cal U}}$ will have to participate in the loss function as a regularization term, see [5], or the approximant has to come from a set with good restricted approximation properties, see Theorem 9.1 and Remark 9.2. In both of these situations there does not appear to exist an optimization procedure with a priori guarantees when using deep neural networks, although using shallow networks a priori bounds can be achieved using a greedy algorithm [66]. In practice, however, once an approximant $v$ is obtained (say, as a result of a minimization of a loss that does not involve $\|v\|_{{\cal U}}$), we can in principle use (7.4) to compute an upper bound for the quality of approximation $\|u-v\|_{H^{1}(\Omega)}$, provided that we can compute an (a posteriori) estimate on $\|v\|_{{\cal U}}$. In fact, similar methods for certifying PINNs solutions have been studied before (see for instance [29] and the references therein).
@@
This error estimator can be used to monitor the error in any optimization scheme that attempts to minimize the loss ${\cal L}^{*}(v)$. For example, suppose that $\Sigma=\Sigma_{n}$ is a linear space of dimension $n$ or a nonlinear manifold determined by $n$ parameters such that each $S\in\Sigma$ is continuous and has a continuous Laplacian $\Delta S$. If at any stage of the optimization procedure we have an $S\in\Sigma_{n}$, then we can use (7.4) to bound the error of $\|u-S\|_{H^{1}(\Omega)}$. Of course, a good estimate requires a bound for $\|S\|_{\cal U}$.
@@
Note that for fixed values of $\widetilde{m}$ and $\overline{m}$, the losses ${\cal L}$ and ${\cal L}^{*}$ are equivalent up to a constant depending upon the parameters $\widetilde{m}$ and $\overline{m}$. This means that if the original loss ${\cal L}$ is driven to 00, i.e. we are perfectly interpolating the data, then both losses give the same error control. However, in the situation where the data is not interpolated, the new loss ${\cal L}^{*}$ gives independent on $m$ control on the error according to Theorem 7.2, while the original PINNs loss ${\cal L}$ does not.
@@
8 Implications for Loss Function Design
In this section, we discuss the implications of Theorem 7.2 on the design of loss functions for the solution of elliptic PDEs using PINNs. Theorem 7.2 implies that the loss function ${\cal L}^{*}$ given in (7.2) gives an a posteriori control on the error up to the optimal recovery rate. However, there are numerous factors that need to be taken into account when designing an appropriate loss function, including the ease of training and implementation. For this reason, directly optimizing the loss ${\cal L}^{*}$ may not be the best choice in practice.
We observe that any loss function ${\cal L}$ which upper bounds ${\cal L}^{*}$ also gives a posteriori error control via Theorem 7.2. With this in mind, we consider the following variants of PINNs for elliptic PDEs. We consider the case $d\geq 3$. Similar discussion holds when $d=2$, taking into account the form of ${\cal L}^{*}$ in (7.2):
  • The original PINNs formulation with an adjustable weight $\lambda>0$ on the boundary terms,
    ${\cal L}_{sq,\lambda}(v):=\left[\frac{1}{\widetilde{m}}\sum_{i=1}^{\widetilde{m}}[\Delta v({\bf x}_{i})+f({\bf x}_{i})]^{2}\right]+\left[\frac{\lambda}{\overline{m}}\sum_{i=1}^{\overline{m}}[v({\bf z}_{i})-g({\bf z}_{i})]^{2}\right].$ (8.1)
    Note that the case $\lambda=1$ corresponds to the original formulation of PINNs [59] and is equivalent to the (square of the) loss ${\cal L}$, see (1.9). Numerous heuristic strategies for choosing the weight $\lambda$ have been explored in a variety of works, see [18, 71, 72].
  • The loss function ${\cal L}^{*}$ in (7.2) with exponent $\tau\geq 1$
    $\begin{split}{\cal L}_{sq,\tau}^{*}(v):=\!\!\left[\frac{1}{\widetilde{m}}\sum_{i=1}^{\widetilde{m}}|f({\bf x}_{i})+\Delta v({\bf x}_{i})|^{\tau}\right]^{2/\tau}\!\!\!\!&+\left[\frac{1}{\overline{m}^{2}}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{\overline{m}}\frac{|[g-v]({\bf z}_{i})-[g-v]({\bf z}_{j})|^{2}}{|{\bf z}_{i}-{\bf z}_{j}|^{2}}\right]\\ &+\left[\frac{1}{\overline{m}}\sum_{j=1}^{\overline{m}}|g({\bf z}_{j})-v({\bf z}_{j})|^{2}\right].\end{split}$ (8.2)
    When $\tau=\gamma$ given in (7.1), this is equivalent to the (square of the) loss ${\cal L}^{*}$.
Note that all of these loss functions are squared, which is typically done to allow for easier optimization (the notation $sq$ here is a shorthand for squared).
We first observe that the $L_{\gamma}$-norm is bounded by the $L_{\tau}$-norm for any $\tau\geq\gamma$ on a finite measure space, such as the bounded domain $\Omega$, or the normalized empirical measure at the data points ${\cal X}$. This means that
$({\cal L}^{*})^{2}\lesssim{\cal L}_{sq,\tau}^{*}\quad\hbox{for any}\quad\tau\geq\gamma,$ (8.3)
and thus Theorem 7.2 gives an a posteriori bound on the solution error in terms of ${\cal L}_{sq,\tau}^{*}$. This means that we can switch to a different norm $L_{\tau}$ with $\tau\geq\gamma$ on the domain term in the loss ${\cal L}^{*}$ and retain an a posteriori error guarantee.
The original PINNs loss ${\cal L}_{sq,\lambda}$ with fixed, independent on $\overline{m}$, $\lambda$ is not an upper bound (with constant independent of $\overline{m}$ or $\widetilde{m}$) on the loss $({\cal L}^{*})^{2}$, and thus does not give an a posteriori error estimate via Theorem 7.2. This is due to the fact that the discrete $L_{2}$-norm is not an upper bound for the discrete $H^{1/2}$-norm on the boundary used in ${\cal L}^{*}$. However, by letting $\lambda$ depend upon the number $\overline{m}$ of the boundary collocation points, one can correct this. Indeed, it follows from (8.3) with $\tau=2>\gamma=\frac{2d}{d+2}$ that
$({\cal L}^{*})^{2}\lesssim{\cal L}^{*}_{sq,2}\lesssim{\cal L}_{sq,\lambda(\overline{m})},\quad\hbox{where}\quad\lambda(\overline{m}):=\overline{m}^{1/(d-1)},$
and the constants in $\lesssim$ are independent of $\overline{m}$. The last inequality is easily proved by observing that
$\frac{1}{\overline{m}^{2}}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{\overline{m}}\frac{|u({\bf z}_{i})-u({\bf z}_{j})|^{2}}{|{\bf z}_{i}-{\bf z}_{j}|^{d}}\leq\frac{2}{\overline{m}^{2}}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{\overline{m}}\frac{|u({\bf z}_{i})|^{2}+|u({\bf z}_{j})|^{2}}{|{\bf z}_{i}-{\bf z}_{j}|^{d}}\lesssim\overline{m}^{-\frac{d-2}{d-1}}\sum_{i=1}^{\overline{m}}|u({\bf z}_{i})|^{2}.$ (8.4)
Remark 8.1.
The value $({\cal L}_{sq,\lambda(\overline{m})}(v))^{1/2}$ can be used in Theorem 7.2 instead of ${\cal L}^{*}(v)$. This indicates how to correctly weight the boundary and domain terms depending on the number of collocation points in the original PINNs formulation, i.e. to use the loss
${\cal L}_{sq,\lambda(\overline{m})}(v)=\left[\frac{1}{\widetilde{m}}\sum_{i=1}^{\widetilde{m}}[\Delta v({\bf x}_{i})+f({\bf x}_{i})]^{2}\right]+\left[\overline{m}^{(2-d)/(d-1)}\sum_{i=1}^{\overline{m}}[v({\bf z}_{i})-g({\bf z}_{i})]^{2}\right].$ (8.5)
All of these alternative loss functions provide an upper bound for the loss ${\cal L}^{*}_{sq,\gamma}$. Consequently, if the (restricted) optimization problem were solved exactly, the loss ${\cal L}^{*}_{sq,\gamma}$ would reach a lower value than these alternatives and would give the best a posteriori error bound up to OR rates. However, in practice, the optimization of neural networks introduces numerous complications, and it is not a priori clear which loss function is best. In §10, we present numerical experiments comparing the practical performance of these loss functions.
Finally, we remark that our analysis of loss functions is performed with respect to the $H^{1}$-error. This same analysis can be performed for different error norms as well (see for instance [78] for the $H^{1/2}$-norm), but we leave this for future work.